Natasha inherits 50,000$ and decides to give 20,000$ to charity. With the remaining money she invest into two stocks paying 16% and 25% . If she makes 6060$ on the two investments, find how much money is invested into each stock.
Answer by checkley77(12844)

50,000-20,000=30,000 for investments.
.25x+.16(30,000-x)=6,060
.25x+4,800-.16x=6,060
.09x=6,060-4,800
.09x=1,260
x=1,260/.09
x=14,000 invested @ 25%.
30,000-14,000=16,000 invested @ 16%.
Proof:
.25*14,000+.16*16,000=6,060
3,500+2,560=6,060
6,060=6,060